### v3

# CodeHalu: Investigating Code Hallucinations in LLMs via Execution-based Verification

Abstract
Large Language Models (LLMs) have made significant progress in code generation, offering developers groundbreaking automated programming support. However, LLMs often generate code that is syntactically correct and even semantically plausible, but may not execute as expected or fulfill specified requirements. This phenomenon of hallucinations in the code domain has not been systematically explored. To advance the community’s understanding and research on this issue, we introduce the concept of code hallucinations and propose a classification method for code hallucination based on execution verification. We categorize code hallucinations into four main types: mapping, naming, resource, and logic hallucinations, with each category further divided into different subcategories to understand and address the unique challenges faced by LLMs in code generation with finer granularity. Additionally, we present a dynamic detection algorithm called CodeHalu designed to detect and quantify code hallucinations. We also introduce the CodeHaluEval benchmark, which includes 8,883 samples from 699 tasks, to systematically and quantitatively evaluate code hallucinations. By evaluating 17 popular LLMs using this benchmark, we reveal significant differences in their accuracy and reliability in code generation, offering detailed insights for further improving the code generation capabilities of LLMs. The CodeHalu benchmark and code are publicly available at https://github.com/yuchen814/CodeHalu.
CodeHalu: Investigating Code Hallucinations in LLMs via
Execution-based Verification
Yuchen Tian1thanks:   Equal contribution. Corresponding to: weixiangyan@ucsb.edu     Weixiang Yan211footnotemark: 1     Qian Yang3,4     Xuandong Zhao5     Qian Chen6 Wen Wang6     Ziyang Luo7     Lei Ma1     Dawn Song5 1The University of Tokyo  2UC, Santa Barbara  3Mila-Québec AI Institute  4Université de Montréal 5UC, Berkeley   6Alibaba Group   7Hong Kong Baptist University yuchentovo@gmail.com      weixiangyan@ucsb.edu      qian.yang@mila.quebec xuandongzhao@berkeley.edu      {tanqing.cq,w.wang}@alibaba-inc.com cszyluo@comp.hkbu.edu.hk      ma.lei@acm.org      dawnsong@cs.berkeley.edu
1 Introduction
Figure 1: The definition and classification of code hallucinations, including 4 main categories and 8 subcategories.
Deep neural networks often generate erroneous information that contradicts the original content, cannot be verified, or conflicts with real-world knowledge. This phenomenon, commonly known as model hallucination, attracts widespread attention in the fields of natural language processing and multimodal learning (Ji et al., 2023; Zhang et al., 2023; Liu et al., 2024), with the community actively exploring methods to mitigate hallucinations (Peng et al., 2023; Elaraby et al., 2023; Liu et al., 2023). However, the issue of model hallucination in the code generation domain remains unexplored.
Conducting a thorough and dedicated study on code hallucinations is crucial for improving the quality of code generated by LLMs. Firstly, the purpose of code is to solve problems, and its value is realized only when the code executes successfully and passes tests (Chen et al., 2021; Austin et al., 2021; Yan et al., 2023). This necessitates that the generated code not only maintain strict logic and precision but also undergoes execution verification to confirm its correctness. Therefore, the practical use and verification of code differ significantly from Natural Language(NL) texts, meaning we cannot directly apply the definitions and methods used for NL hallucinations to code. Secondly, code snippets containing hallucinations may trigger runtime errors, or exhibit functional defects, which hinder the reliable deployment of LLMs in automated software development scenarios. Lastly, by exploring and verifying code hallucinations in a targeted manner, we can effectively uncover their causes and contribute to improving the architecture and training methods of LLMs.
To fill this gap, we define the concept of code hallucination in LLMs, based on the unique purpose and function of the code. Code hallucinations refer to the phenomenon where code generated by LLMs is syntactically correct or even semantically plausible but ultimately cannot execute as expected or fails to meet specified requirements. [^1] We test 16 LLMs using 105,958 code samples. The experimental results demonstrate that only 9 models occasionally exhibit syntactic errors in the generated code, with an exceptionally low average error rate of 0.0020. These findings support our initial hypothesis that the code generated by LLMs is generally syntactically correct and even semantically plausible or appropriate. Detailed statistical data and experimental results are shown in Appendix A. This phenomenon typically arises from various factors, such as errors or outdated information in the training data, an inadequate grasp of the syntax rules and programming paradigms of the programming languages, and limitations in the logical processing capabilities of the models. In contrast to previous methods that passively explore hallucinations in NLP through a Q&A framework or by prompting LLMs to generate hallucinated answers (Lin et al., 2021; Cheng et al., 2023), we employ an active strategy to detect hallucinations during the code generation process by LLMs. This approach is crucial as the ultimate goal of the generated code is to execute correctly and fulfill specific tasks.
To detect and quantify hallucinations in LLMs during code generation, we develop a dynamic detection algorithm named CodeHalu. This algorithm employs a statistical induction method based on execution validation to identify specific patterns that frequently occur in code generated by multiple LLMs, such as error types, syntax interruptions, or unexpected execution results. When a pattern consistently appears across multiple LLMs, it is recognized as a common code hallucination. Based on the CodeHalu algorithm 1, we employ an execution-based validation approach for hallucination detection, combined with a two-stage heuristic identification method. By conducting statistical quantification on 17 mainstream LLMs, we categorize code hallucinations into four major categories: Mapping, Naming, Resource, and Logical Hallucinations. These categories are further divided into eight subcategories, as illustrated in Figure 1. We analyze 17 LLMs for cross-task occurrence rates in eight categories of code hallucinations. The low average rate of 2.04% confirms the independence and validity of our classification.
To effectively measure and compare code hallucinations across different LLMs, we introduce an evaluation benchmark named CodeHaluEval, which is based on the incidence rate of hallucinations. It follows a structured process of Validation-Identification-Construction as shown in Figure 4 to detect and evaluate code hallucinations in LLMs, closely tied to real-world programming scenarios, ensuring that the generated code correctly achieves the expected functionality. CodeHaluEval encompasses eight types of code hallucinations as illustrated in Figure 1, covering 699 distinct tasks and corresponding to 8,883 samples. Additionally, we systematically evaluate 17 mainstream LLMs to reveal the distribution and behavior patterns of their code hallucinations. We also analyze the potential causes of various code hallucinations, providing detailed insights for further improving the code generation capabilities of LLMs. Our contributions can be summarized as follows:
  • Code Hallucination: We introduce the concept of code hallucination in LLMs and propose an execution-based verification method to define code hallucination, addressing a gap in the research on hallucination within the code generation domain.
  • CodeHalu Algorithm: We develop a dynamic detection algorithm, CodeHalu, to identify and quantify the types of hallucinations that occur in LLMs during code generation. We categorize code hallucinations into four main categories based on a two-stage heuristic approach, discussing their theoretical implications and potential causes.
  • CodeHaluEval Benchmark: We propose the CodeHaluEval benchmark to systematically evaluate 17 popular LLMs, revealing the distribution and patterns of code hallucinations across these models, and providing insights for developing more robust and reliable LLMs.
2 Related Work
Hallucination
In the field of NLP, hallucination is initially defined as the phenomenon where the text generated by a model is fluent and natural but either lacks substantive meaning or is inconsistent with the provided source content (Ji et al., 2023). Recently, Zhang et al. (2023) standardize the definition of NL hallucinations in LLMs into three categories: input-conflicting hallucinations, where the content generated by LLMs diverges from the user’s input; context-conflicting hallucinations, in which the generated content contradicts previously generated content; and fact-conflicting hallucinations, where the generated content conflicts with established world knowledge. These hallucinations are attributed to various factors, such as poor-quality data samples in the training dataset or the use of sampling algorithms with high uncertainty.
In the multimodal domain, Zhai et al. (2023) classify types of hallucinations in image-to-text scenarios, such as image captioning and visual question answering. They define three main types of hallucinations: object existence hallucinations, object attribute hallucinations, and object relationship hallucinations. In text-to-image scenarios, such as image generation, hallucinations refer to the creation of factually incorrect details by the image generation model in response to the given text input. Huang et al. (2024) introduce VHTest, which evaluates hallucinations across eight dimensions in images, including the existence, shape, color, orientation, OCR, size, position, and counting of visual objects. In text-to-video scenarios, such as video generation, Chu et al. (2024) define three types of hallucinations: prompt consistency hallucinations, static hallucinations, and dynamic hallucinations. Detailed definitions and descriptions of each type of hallucination are provided in Appendix A. Although the issue of hallucinations receives extensive attention in NLP and multimodal domains, it remains unexplored in the code domain. Therefore, we propose CodeHalu to systematically define, identify, classify, and quantify code hallucinations in LLMs.
Existing Coding Benchmarks
In recent years, numerous studies focus on evaluating the capability of LLMs to handle various programming tasks. Among these, the HumanEval (Chen et al., 2021), includes 164 Python programming problems, each with an average of 6.7 unit tests. The MBPP (Austin et al., 2021) benchmark contains 974 Python programming tasks. Compared to these, the APPS (Hendrycks et al., 2021) benchmark presents more challenging programming questions, with each problem averaging 293.2 words in length. CodeScope (Yan et al., 2023) covers 43 programming languages and eight coding tasks to comprehensively evaluate LLMs in code understanding and generation. MMCode (Li et al., 2024) is designed to evaluate the programming capability of code models in multimodal scenarios. SWE-bench (Jimenez et al., 2023) evaluates the capability of LLMs to edit code repositories to solve problems with a level of complexity similar to that faced by human programmers in real-world programming situations. Overall, existing code benchmarks focus on evaluating the performance of LLMs on various programming tasks. However, there is still a lack of effective methods to detect and quantify potential hallucinations that may occur in code generation. Therefore, we propose CodeHaluEval to detect and quantify code hallucinations in LLMs.
3 Code Hallucination
As a tool, code aims to achieve specific objectives through correct execution. This inherent characteristic motivates our use of an execution-based verification method to explore and identify code hallucinations. In this section, we define the concept of code hallucination and distinguish it from code errors, clarifying the relationship and differences between these two phenomena.
Definition 1 (Code Hallucinations).
Code hallucinations refer to the code generated by large language models that is syntactically correct or even semantically plausible, but ultimately cannot execute as expected or fails to meet specified requirements.
Definition 2 (Code Errors).
Code errors refer to issues in a program that cause it to stop executing.
Remark 3 (Code Hallucinations vs. Code Errors).
In multiple domains, existing work (Ji et al., 2023; Zhang et al., 2023; Huang et al., 2024; Zhai et al., 2023; Chu et al., 2024) often equates errors with hallucinations, or considers errors as a specific subset of hallucinations. We follow this perspective and regard code errors as a specific subset of code hallucinations. In other words, errors manifest as a form of hallucination, but not all hallucinations can be adequately expressed through errors. Figure 2 illustrates the distinction between code errors and code hallucinations. The code on the left exhibits a typical code error due to the use of an undefined variable “N”, resulting in a NameError. On the right, the code repeatedly calls the same function due to a logical collapse during generation, eventually exceeding the maximum token limit and leading to a SyntaxError. However, the underlying issue is a latent logical hallucination, rather than the observed syntactic error.
Overall, although there is some slightly overlap between code hallucinations and code errors, their meanings, research objects, and scopes differ significantly. Code hallucinations focus on why the model produces hallucinations, while code errors focus on what grammatical rules the code violates. Code errors form a proper subset of code hallucinations, while code hallucinations encompass a broader range of potential logical and functional issues, representing a finer-grained and more comprehensive evaluation of the overall quality and functionality of the code.
Figure 2: Examples that differentiate between code errors and code hallucinations.
4 CodeHalu Algorithm
Algorithm 1 CodeHalu Algorithm
1:Input: Code Generation Dataset $\alpha$, Language models $\mathsf{\pi}$
2:Output: HaluTypes $\xi$
3:Let $\xi\leftarrow$ empty list
4:for $\alpha_{i}$, where $i\in\{1,\dots,k\}$ do
5:    for $\pi_{j}$, where $j\in\{1,\dots,m\}$ do
6:         $\mathsf{GC}_{j}^{\alpha_{i}}\leftarrow$ $\pi_{j}(\mathsf{GI}_{j},\mathsf{Q})$
7:         if $\mathsf{GC}_{j}^{\alpha_{i}}$ is stuttering, infinite enumeration, or gibberish
8:             $\xi\leftarrow\xi\cup\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$
9:         else
10:             for $t_{n}$, where $n\in\{1,\dots,N\}$ do
[…]
12:                         $\mathsf{ER}_{j}^{\alpha_{i}}(t_{n})\leftarrow$Execute$(\mathsf{GC}_{j}^{\alpha_{i}}(t_{n}))$
13:                         if $\mathsf{ER}_{j}^{\alpha_{i}}(t_{n})\neq op_{t_{n}}$
14:                             $\xi\leftarrow\xi\cup\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$
15:                 else
16:                         $\xi\leftarrow\xi\cup\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$
17:Aggregate and count frequencies of unique $\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$ in $\xi$
In this section, we introduce a dynamic detection algorithm called CodeHalu, which detects and quantifies hallucinations in LLMs in code generation. CodeHalu operates on the assumption $\mathsf{ASS}$: if a specific pattern frequently appears in the code generated by multiple LLMs, it is considered a common code hallucination. These patterns include error types, syntax interruptions, logical collapse, or unexpected execution results.
Consider a dataset $\alpha$ contains $k$ samples, where each sample $\alpha_{i}$ consists of a problem description $Q$ and a series of test cases $t_{1},t_{2},\ldots,t_{n}$. Each test case $t_{n}$ includes an input $\mathsf{ip}$ and the corresponding expected output $\mathsf{op}$. Notably, following previous work (Li et al., 2023b; Yan et al., 2023), we integrate resource (time and memory) constraints into the code generation instructions. As shown in Algorithm 1, we use a $\mathsf{\pi}_{j}$ to generate a code solution $\mathsf{GC}_{j}^{\alpha_{i}}$ for each sample $\alpha_{i}$ based on the code generation instruction $\mathsf{GI_{j}}$ and the problem description $\mathsf{Q}$. If $\mathsf{GC}_{j}^{\alpha_{i}}$ exhibits any of the states such as stuttering, infinite loops, or gibberish, we include it in $\xi$.
To test the potential hallucinations of $\mathsf{GC}_{j}^{\alpha_{i}}$ at a fine-grained level, we execute all test cases of sample $\alpha_{i}$ one by one to verify whether it successfully executes and meets the expected functionality. We record the actual execution result $\mathsf{ER}_{j}^{\alpha_{i}}(t_{n})$ of the code under each test case $t_{n}$ and extensively test each sample $\alpha_{i}$ across more than 15 $\mathsf{\pi}$ to obtain statistically-based inductive results. If the code execution fails or does not meet the expected results, we record it in $\xi$.
Finally, we merge the identical states $\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$ detected by CodeHalu and calculate their occurrence frequencies. According to assumption $\mathsf{ASS}$, $\xi$ can be represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{o}$, $P_{o}$)], where $\xi_{o}$ denotes the $o^{th}$ type of code hallucination, and $P_{o}$ indicates its frequency. Code hallucinations come from four perspectives: errors, syntax, logic, and execution results. Additionally, CodeHalu is language-agnostic and can adjust to match various programming scenarios depending on the programming language.
5 Code Hallucinations Classification
In this section, we analyze the hallucination states detected by the CodeHalu algorithm, classify and define four main types of hallucinations, and discuss the rationale behind the classification method.
According to the TIOBE Index [^2] https://www.tiobe.com/tiobe-index/, a metric of programming language popularity, we primarily investigate code hallucinations within the Python. By applying the CodeHalu algorithm on the complex APPS dataset (Hendrycks et al., 2021) and 17 widely-used LLMs, we identify and validate 18 types of hallucination states that violate human expectations during the code generation, including inconsistent code context, ambiguous logic and data flow, conflicting intentions, among others. Using the two-stage heuristic classification method introduced in Remark 8, we categorize code hallucinations into four main types based on the nature and origin of these phenomena: mapping hallucinations, naming hallucinations, resource hallucinations, and logical hallucinations, as illustrated in Figure 1.
Definition 4 (Mapping Hallucinations).
Mapping Hallucinations refer to the ambiguity and confusion that occur in LLMs’ perception and mapping of data types, values, and structures during data operations. This phenomenon is further divided into two sub-categories: data compliance hallucinations and structure access hallucinations.
Data compliance hallucinations occur when LLMs have a vague understanding of the data types and parameter values of the objects being manipulated, resulting in generated code that attempts to perform type-mismatched or rule-violating operations.
Structure access hallucinations occur when LLMs misinterpret the data structures of the objects being manipulated, leading to generated code that attempts to access non-existent array indices or dictionary keys.
Definition 5 (Naming Hallucinations).
Naming Hallucinations refer to the memory-related issues and factual inaccuracies exhibited by LLMs when handling the naming, scope, and existence of variables, attributes, and modules. This phenomenon is further divided into two subcategories: identity hallucinations and external source hallucinations.
Identity hallucinations occur when LLMs possess biased memories or lack sufficient understanding of the context, leading to generated code that references undefined variables, accesses non-existent object properties, or uses unassigned variables in local scopes.
External source hallucinations occur when LLMs exhibit significant memory-related issues or obvious conflicts with facts concerning external knowledge sources, resulting in generated code that attempts to import non-existent modules or fails to correctly load modules from other paths.
Definition 6 (Resource Hallucinations).
Resource Hallucinations occur when LLMs lack an adequate perception and prediction of resource consumption and control flow of the generated code during execution. This phenomenon is further divided into physical constraint hallucinations and computational boundary hallucinations.
Physical constraint hallucinations arise when LLMs underestimate resource consumption during data processing operations, causing code failure due to exceeding memory capacity, stack depth, or other physical constraints.
Computational boundary hallucinations occur when LLMs blur recognition of numerical calculation limits and iteration endpoints during data processing operations, causing code failure due to numerical overflow or improper iteration control.
Definition 7 (Logic Hallucinations).
Logic Hallucinations refer to the discrepancies between the expected results and the actual outcomes after executing the code generated by LLMs, or outputs with low semantic density or even complete chaos. This phenomenon is further divided into logic deviation and logic breakdown.
Logic deviation occurs when LLMs generate code that lacks sufficient logical consideration or contradicts the intended instructions. While this hallucination may not cause errors during execution, logical deviations or confusion result in outcomes that fail to meet the expected results.
Logic breakdown occurs when LLMs struggle to interpret or maintain a continuous understanding of context during code generation. This indicates that the models may lose direction while generating code, making it difficult to maintain strict consistency of contextual information.
Remark 8 (Discussion of Rationality).
To ensure the rationality and effectiveness of our code hallucination classification method, we conduct in-depth analyses.
Firstly, we extensively reference classification methods for hallucinations in the fields of NLP and multimodal research (Zhang et al., 2023; Ji et al., 2023; Huang et al., 2024; Zhai et al., 2023; Chu et al., 2024), as well as methods for classifying code errors and vulnerabilities in software engineering (Pan et al., 2023; Wang et al., 2024; Huang et al., 2023). We adopt a two-stage heuristic classification strategy. Initially, team members independently review failed code cases and develop preliminary classification frameworks; then, we reach a consensus through collaborative discussions. This widely used approach ensures the adaptability and accuracy of our framework, enabling a systematic understanding of code hallucinations in LLMs.
Secondly, we analyze the cross-task occurrence rates of each model across eight categories of code hallucinations, detailed in Appendix Table 3. The results show that the average cross-task occurrence rate for these categories is only 2.04%, confirming the independence and rationality of our classification. For the Gemma-7B model, which exhibits the most severe hallucinations in Table 2, only 1.07% of task samples show cross-task hallucinations, as illustrated in Figure 3.
Lastly, we conduct an empirical investigation of our classification results and design a questionnaire to evaluate the rationality of our method, detailed in the Appendix Figure 8. The survey receives 23 responses, and after excluding seven respondents with less than three years of experience, we analyzed 16 valid responses. The survey results indicate a rationality rating of 91.08% for our classification method, further supporting its validity.
Figure 3: The diagram illustrates the intersection of various hallucinations in Gemma-7B during the CodeHaluEval. The bar chart at the top shows the frequency of each intersection, while the bar chart on the left indicates the frequency of each type of hallucination. The connecting lines represent the co-occurrence patterns between different hallucinations.
Figure 4: Collection of CodeHaluEval benchmark based on a verification-identification-construction process.
6 Cause Analysis of Code Hallucinations
In this section, we explore the potential causes of various hallucinations generated by LLMs, aiming to provide valuable insights for optimizing training data, training methods, model architecture, and alignment strategies.
Mapping hallucinations stem from the model’s misunderstanding of data types and structures. This phenomenon arises due to several factors: (1) The model generates code based on tokens, lacking insight into higher-level structures such as statements and functions (Yang et al., 2021); (2) When dealing with long-distance data dependencies, especially within complex code blocks, the model fails to continuously track the structure and state of variables, overly relying on local information while neglecting the importance of the overall context (Zhang et al., 2024); (3) The model does not explicitly perform type checking and structure matching during code generation, lacking static checking and error correction mechanisms.
Naming hallucinations reflect the limitations of models in tracking information and utilizing external knowledge. This issue arises from several factors: (1) Token-based feature representation makes it difficult to accurately model long-distance dependencies, leading to model misjudgments regarding variable scope, lifecycle, and visibility (Xu et al., 2020); (2) The code generation process lacks consistency checks for identifiers and does not perform global tracking of variable definitions and usage; (3) Knowledge of external libraries is not effectively and timely integrated into the model’s knowledge system, making it difficult for the model to accurately understand the names, functions, and invocation methods of libraries (Jesse et al., 2023).
Resource hallucinations highlight the model’s lack of deep understanding of code execution mechanisms and physical constraints. These issues arise from several factors: (1) The training data lacks information related to resource consumption and performance optimization, making it difficult for the model to learn about complexity analysis and resource evaluation; (2) As the model generates code based on probabilities, it lacks a module for calculating and estimating the resource consumption of the generated code, making it unable to simulate the real-world operating environment and resource limits; (3) During the model training process, the focus is on the correctness of the code’s functionality, often overlooking its complexity and resource constraints in actual execution environments.
Logic hallucinations reveal the model’s deficiencies in semantic understanding and reasoning about code. This issue arises due to several factors: (1) The model relies on pattern matching and statistical rules to generate code, lacking a fundamental understanding of symbolic systems and rigorous verification of program logic; (2) The training data is often not rigorously verified for accuracy and may contain code with very similar functions. Since models sometimes imitate and memorize previous examples (Yan and Li, 2022), this can result in the model directly replicating similar logic in the code or even learning incorrect logic from the outset; (3) When the model generates code, repetition at the line level has a self-reinforcing effect, causing the model to become increasingly confident in the code it generates, which may lead to a stuttering phenomenon (Xu et al., 2022).
7 The CodeHaluEval Benchmark
We construct the CodeHaluEval benchmark, a unified evaluation method for comparing various types and frequencies of hallucinations in code generation across different LLMs. We develop the CodeHaluEval based on the APPS testing set, following a structured process of Validation-Identification-Construction, as shown in Figure 4.
Category #Tasks #Samples Sub-Category #Tasks #Samples Mapping 262 2,288 Data Compliance 110 941 Structure Access 152 1347 Naming 157 1,853 Identity 115 1323 External Source 42 530 Resource 107 1,130 Physical Constraint 47 491 Computational Boundary 60 639 Logic 173 3,612 Logic Deviation 119 2,443 Logic Breakdown 54 1,169
Table 1: Detailed statistics of categories, and quantities in CodeHaluEval benchmark.
Model Mapping ($\downarrow$) Naming ($\downarrow$) Resource ($\downarrow$) Logic ($\downarrow$) Average ($\downarrow$) DC SA Avg. ID ES Avg. PC CB Avg. LD LB Avg. GPT-4 32.31 10.02 19.19 27.74 0.57 19.97 0.20 3.76 2.21 85.76 0.51 58.17 33.04 LLaMA-3-8B 46.87 23.46 33.09 12.09 0.00 8.63 15.48 12.99 14.07 78.39 0.00 53.02 33.67 DeepSeek Coder-6.7B 24.23 25.61 25.04 15.80 0.00 11.28 17.52 17.21 17.35 99.06 0.17 67.05 38.28 GPT-3.5 20.19 22.05 21.28 30.54 0.00 21.80 18.53 6.42 11.68 99.88 0.00 67.55 38.98 Claude-3-haiku 38.68 29.25 33.13 9.07 0.75 6.69 37.07 20.81 27.88 100.00 0.17 67.69 41.00 ChatGLM-3-6B 36.13 44.91 41.30 50.87 0.00 36.32 24.85 2.35 12.12 88.99 0.00 60.19 44.23 Ernie-3.5 48.14 36.90 41.52 30.31 0.38 21.75 18.13 11.89 14.60 98.98 0.00 66.94 44.31 Qwen-turbo 49.63 48.33 48.86 29.48 2.08 21.64 7.94 2.82 5.04 98.08 0.17 66.39 44.74 MagicCoder-7B 50.27 26.58 36.32 17.69 0.00 12.63 21.18 28.33 25.22 100.00 16.25 72.90 44.84 Code LLaMA-7B 65.04 42.17 51.57 31.07 0.00 22.18 18.53 6.26 11.59 94.76 9.41 67.14 46.68 StarCoder-16B 48.14 38.83 42.66 60.70 9.25 45.98 28.92 11.11 18.85 95.09 0.77 64.56 49.23 LLaMA-2-7B 51.22 32.29 40.08 78.46 71.13 76.36 14.87 0.00 6.46 81.05 0.34 54.93 49.41 Gemini-1.0 34.11 53.53 45.54 45.88 0.00 32.76 24.44 16.12 19.73 98.65 10.35 70.07 49.57 Mistral-7B 45.48 36.53 40.21 59.18 15.85 46.79 27.49 10.80 18.05 99.35 0.17 67.25 49.76 WizardCoder-7B 26.57 31.40 29.41 31.29 0.00 22.34 33.20 9.39 19.73 93.90 72.37 86.93 50.10 CodeGeeX-2-6B 47.61 27.99 36.06 45.05 0.00 32.16 36.66 23.47 29.20 89.60 99.66 92.86 57.47 Gemma-7B 55.26 41.05 46.90 51.85 0.00 37.02 14.46 14.55 14.51 97.18 100.00 98.09 61.53
Table 2: Evaluation results of 17 models on CodeHalu. DC denotes Data Compliance hallucination. SA denotes Structure Access hallucination. ID denotes identity hallucination. ES denotes External Source hallucination. PC denotes Physical Constraint hallucination. CB denotes computational Boundary hallucination. LD denotes Logic Deviation. LB denotes Logic Breakdown.
In the validation phase, we use the CodeHalu algorithm to identify multiple types of hallucinations $\mathsf{HaluTypes}$ $\xi$, represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{i}$, $P_{i}$)]. In the identification phase, we annotate the ${3k}$ most common hallucinations and their frequencies in each sample $\alpha_{i}$, represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{3k}$, $P_{3k}$)]. In the construction phase, we sort all samples in descending order based on the frequency $P_{i}$ of each hallucination type $\xi_{i}$. If the hallucination frequency in a sample $\alpha_{i}$ exceeds the threshold $k$, we include this sample in the corresponding hallucination type set in the CodeHaluEval benchmark. When selecting the threshold $k$, we consider both the minimum number of samples required to detect code hallucination effects in the CodeHaluEval benchmark and the inference costs associated with evaluating various LLMs. The efficacy analysis and the selection process for the threshold $k$ are detailed in Appendix A. Through this method, we establish the CodeHaluEval benchmark, with detailed statistics shown in Table 1.
8 Experiments
Models. To comprehensively analyze the different hallucinations of various competitive LLMs in CodeHaluEval, we evaluate 12 general LLMs, including GPT-4 (OpenAI, 2023), GPT-3.5 (OpenAI, 2023), Gemini-Pro-1.0 (Gemini, 2023), Claude-3-haiku (Anthropic, 2024), LLaMA-2 & 3 (Touvron et al., 2023), Vicuna (Chiang et al., 2023), Qwen-turbo (Bai et al., 2023), ChatGLM3-6B (Du et al., 2021), Ernie-3.5(Baidu, 2023), Mistral-7B(Jiang et al., 2023), Gemma (Team et al., 2024). We also evaluate 5 coding LLMs, including Code LLaMA (Roziere et al., 2023), DeepSeek Coder (Guo et al., 2024), CodeGeeX-2 (Zheng et al., 2023), StarCoder-2 (Li et al., 2023a), MagicCoder-7B (Wei et al., 2023), WizardCoder-7B (Luo et al., 2023). The experimental evaluation is conducted using API calls or 8 NVIDIA A6000 GPUs.
Metrics. Given the limited exploration of code hallucinations, no dedicated metrics currently exist for evaluating them in LLMs. To address this gap, we propose an evaluation metric called Hallucination Rate (HR). Specifically, HR is defined as the percentage of hallucination samples detected in the test set among all samples, with the formula: $\operatorname{HR}=\frac{1}{N}\sum_{i=1}^{N}S(i,K)$, where $S(i,K)$ is an indicator function. If the $i^{th}$ sample satisfies the hallucination condition, then $S(i,K)$ = 1; otherwise, $S(i,K)$ = 0. Ideally, a lower HR indicates a lower likelihood of hallucinations during code generation by the LLM, thus demonstrating greater robustness and reliability. To our knowledge, HR is the first metric that accurately reflects the hallucination phenomenon in LLMs during code generation tasks through actual execution tests.
Result & Analysis
Figure 5: The performance of 17 LLMs on different types of hallucinations and the overall hallucination rate.
The experimental results are presented in Table 2 and Figure 5, with case studies of various models across different hallucination categories detailed in Appendix A.
Mapping hallucination: GPT-4 and GPT-3.5 consistently identify and follow rules related to data types, values, and structures, demonstrating strong context sensitivity.
Naming hallucination: Claude-3 reliably remembers and references entity names from the context and external knowledge bases. In contrast, LLaMA-2 exhibits significant memory bias when processing external knowledge and occasionally fabricates information.
Resource hallucination: GPT-4, Qwen, and LLaMA-2 effectively account for actual resource constraints when generating code, showing an understanding of computational boundaries and limitations, which leads them to produce code with lower complexity.
Logical hallucination: Although all models face challenges in maintaining logical coherence, LLaMA-3 and GPT-4 perform relatively well in reducing repetition. Most models rarely generate code with stuttering or infinite loops, but such issues are more common in Gemma, CodeGeeX-2, and WizardCoder, indicating a tendency to lose semantic and logical consistency during code generation.
Overall, GPT-4 and LLaMA-3 perform well across all hallucination categories, displaying stability and robustness in various scenarios. Logical hallucinations remain the most prevalent issue across all models, while naming and resource hallucinations are relatively less common. The performance of different models varies significantly across hallucination types, likely due to differences in their training data, methods, and architectures. The average hallucination rate ranges from approximately 20% to 60%.
We view mitigating code hallucination as future work. Based on a detailed analysis of experimental results and generated cases, we provide insights into strategies for mitigating code hallucinations in LLMs. In terms of training data, improving the quality and increasing the diversity of data sources enhances the model’s generalization ability. In terms of training methods, employing alignment strategies based on compilation and execution verification, as well as setting multiple objectives during training, enables the model to better understand the data flow and control flow of code. In terms of model architecture, introducing a static code verification module provides real-time feedback on verification results, thereby enhancing the model’s robustness. Additionally, incorporating a code graph module allows the model to construct and utilize graph structure information when generating code, deepening its understanding of patterns and logical relationships in the generated code.
9 Conclusion
We introduce the concept of code hallucination and propose an execution-based verification method to classify code hallucinations. We develop the dynamic detection algorithm, CodeHalu, which categorizes code hallucinations into four main types and eight subtypes, providing a comprehensive understanding of the various challenges faced by LLMs in code generation. Additionally, we establish the CodeHaluEval benchmark and evaluate 17 widely-used LLMs, revealing significant differences in their hallucination patterns during code generation, and providing detailed insights for further improving the code generation capabilities of LLMs. Overall, we lay the theoretical foundation for understanding the hallucination phenomenon of LLMs in code generation, and provide a complete set of tools for detecting and evaluating code hallucinations.
10 Limitations
Python is our focus for exploring code hallucination, as it is the most widely used programming language according to the TIOBE Index. Furthermore, many existing studies, such as HumanEval and MBPP benchmarks, concentrate on Python. Thus, we do not extend our investigation to other languages.
CodeHalu focuses on ensuring the correctness of generated code to meet the needs of developers and users. In contrast, identifying and preventing security risks is a higher-level concern, effectively addressed through sandbox environments.
We focus on code hallucination specifically within the code generation task, excluding other programming tasks such as code translation, and code repair. This is because code generation is currently the most widely studied task in the community. It is important to emphasize that our hallucination detection and evaluation methods can be easily adapted to other tasks, which we consider as future work.
Appendix A Appendix
SyntaxError Analysis
Figure 6: The frequency of syntactic errors in the code generated by various models. The frequency of errors equals the number of erroneous samples divided by the total number of samples, rounded to six decimal places.
Related Work
In the field of NLP, hallucination is initially defined as the phenomenon where the text generated by the model is fluent and natural but lacks actual meaning or is inconsistent with the provided source content (Ji et al., 2023). Recently, Zhang et al. (2023) standardize the definition of NL hallucinations in LLMs into three categories: Input-conflicting hallucinations, where the content generated by LLMs diverges from the user’s input; Context-conflicting hallucinations, in which the generated content contradicts previously generated content; and Fact-conflicting hallucinations, where the generated content conflicts with established world knowledge. These hallucinations may be attributed to various factors, such as poor-quality data samples in the training dataset or the use of sampling algorithms with high uncertainty. To mitigate the impact of these hallucinations, recent studies primarily focus on addressing the issues of data preparation phase (Gardent et al., 2017), training phase (Elaraby et al., 2023), and inference phase (Dhuliawala et al., 2023) to alleviate the hallucination problem in LLMs.
In the multimodal domain, Zhai et al. (2023) classify types of hallucinations in image-to-text scenarios, such as image captioning and visual question answering. They define three main types of hallucinations: object existence hallucinations, where an object mentioned in the description does not exist in the corresponding image; object attribute hallucinations, where the description of the object’s attributes does not match the actual attributes in the image; and object relationship hallucinations, where the relationships between objects are incorrectly described. In text-to-video scenarios, such as video generation, Chu et al. (2024) define three types of hallucinations: prompt consistency hallucinations, which means the text description does not match the visual output; static hallucinations, which refers to errors in the spatial relationships, physical properties, and semantic consistency of objects and scenes within a single frame; and dynamic hallucinations, which indicates inconsistencies or abnormalities in the motion and behavior of objects or entities across frames.
CodeHaluEval Power Analysis
To ensure our study has sufficient statistical power to detect hallucinations commonly triggered by different LLMs during code generation tasks, we conduct a power analysis to determine the minimum sample size required. The power analysis is based on the following parameters, which are standard in social and behavioral science research:
  • Power Level: We set the power level at 0.8 to ensure adequate statistical strength for detecting actual effects.
  • Significance Level (Alpha): We set the significance level at 0.05 to balance the study’s sensitivity with the false positive rate.
  • Effect Size: We select a Cohen’s d value of 0.2, a widely accepted standard for small effect sizes, to ensure reliable detection of even relatively minor effects.
Using the statsmodels library for calculations, we determine that a minimum of 394 samples is necessary. This indicates that at least 394 code task samples are required to ensure the statistical reliability of our research findings.
In selecting the $k$-value (defined as the minimum number of models that must identify the same hallucination for a task to be classified as prone to triggering hallucinations), we perform the following analysis:
As illustrated in Figure 7, we summarize the number of samples included in CodeHaluEval for different $k$-values. At $k$ = 5, we find that 699 tasks are identified by at least five models as potentially triggering hallucinations, which far exceeds our minimum sample size of 394 derived from power analysis. This ensures the reliability and robustness of our statistical analysis.
If the $k$-value is set higher than 5, the number of tasks identified as likely to trigger hallucinations falls below 394, leading to insufficient statistical power and increasing the risk of unstable research results.
In conclusion, we choose $k$ = 5 as the threshold, not only meeting the requirements for statistical power but also increasing the sensitivity to accurately identify true hallucination scenarios while maintaining a lower misidentification rate. This methodological choice provides strong statistical support for our benchmark tests and ensures the reliability and practicality of our results.
Figure 7: Number of samples retained at different threshold levels k. When k reaches 6, the sample size is 331, which does not meet the minimum requirement of 394 samples.
Classification of Code Hallucinations
Model #2($\downarrow$) #3($\downarrow$) Avg.($\downarrow$)
LLaMA-2-7B 0.72 0.00 0.36
Gemma-7B 2.15 0.00 1.07
ChatGLM-3-6B 2.15 0.14 1.14
GPT-4 2.86 0.00 1.43
GPT-3.5 3.00 0.14 1.57
Claude-3-haiku 3.15 0.00 1.57
DeepSeek Coder-6.7B 3.29 0.00 1.65
MagicCoder-7B 3.29 0.14 1.72
WizardCoder-7B 3.58 0.00 1.79
Gemini-1.0 3.86 0.29 2.07
Code LLaMA-7B 4.43 0.00 2.22
Mistral-7B 4.43 0.14 2.29
Qwen-turbo 4.86 0.00 2.43
StarCoder-16B 5.01 0.29 2.65
Ernie-3.5 5.29 0.43 2.86
CodeGeeX-2-6B 11.44 0.14 5.79
[…]
Table 3: Task-level hallucination crossover rates. #2 denotes the scenario where two types of hallucinations are detected simultaneously in a single task, and #3 denotes the scenario where three types of hallucinations are detected simultaneously in a single task. Lower crossover rates indicate that the hallucinations are independent of each other.
Data compliance hallucinations occur when LLMs have a vague understanding of the data types and parameter values of the objects being manipulated, resulting in generated code that attempts to perform type-mismatched or rule-violating operations. This hallucination typically manifests as TypeError, ValueError, or ZeroDivisionError exceptions being thrown during code execution, reflecting the model’s unexpected behavior in data type validation, parameter value handling, and arithmetic operations during code generation.
Structure access hallucinations occur when LLMs misinterpret the data structures of the objects being manipulated, leading to generated code that attempts to access non-existent array indices or dictionary keys. This hallucination typically manifests as IndexError or KeyError exceptions being thrown during code execution, reflecting the model’s lack of clarity in understanding the data structures within the given context.
Identity hallucinations occur when LLMs possess biased memories or lack sufficient understanding of the context. This leads to generated code that references undefined variables, accesses non-existent object properties, or uses unassigned variables in local scopes. This hallucination typically manifests as NameError, AttributeError, or UnboundLocalError exceptions being thrown during code execution, reflecting the model’s difficulty in managing long-distance dependencies and accurately tracking variable definitions and scopes.
External source hallucinations occur when LLMs exhibit significant memory-related issues or obvious conflicts with facts concerning external knowledge sources, resulting in generated code that attempts to import non-existent modules or fails to correctly load modules from other paths. This hallucination typically manifests as ImportError or ModuleNotFoundError exceptions being thrown during code execution, reflecting the model’s lack of deep understanding and accurate memory of the representation and organization of external knowledge sources.
Physical constraint hallucinations occur when LLMs underestimate resource consumption during data processing operations, causing code failure during execution due to exceeding memory capacity, stack depth, or other physical constraints. This hallucination typically manifests as RecursionError or MemoryError exceptions being thrown during code execution, reflecting the model’s inadequate understanding of the global operating conditions.
Computational boundary hallucinations occur when LLMs blur recognition of numerical calculation limits and iteration endpoints during data processing operations, causing code failure due to numerical overflow or improper iteration control. This hallucination typically manifests as OverflowError or StopIteration exceptions being thrown during code execution, reflecting the model’s erroneous understanding of numerical boundary conditions and insufficient grasp of the inherent logic of iterators, loops, and control flows.
Logic deviation occurs when LLMs generate code that lacks sufficient logical consideration or contradicts the intended instructions. This hallucination usually does not cause errors during code execution; however, due to logical deviations or confusion, the program’s outcomes fail to meet the expected results. This issue reflects the model’s insufficient handling of logic and inadequate understanding of algorithms during the code generation process.
Logic breakdown occurs when LLMs struggle to interpret or maintain a continuous understanding of context during code generation. This indicates that the models may lose direction while generating code, making it difficult to maintain strict consistency of contextual information. Consequently, the models might generate stuttering, infinite enumeration, or gibberish code, often lacks structural coherence and logical consistency.
Case Study
Table 4-A provide case studies of different models in various types of code hallucinations.
NL-Description
Salem gave you $n$ sticks with integer positive lengths $a_{1},a_{2},\ldots,a_{n}$. For every stick, you can change its length to any other positive integer length (that is, either shrink or stretch it). The cost of changing the stick’s length from $a$ to $b$ is $|a-b|$, where $|x|$ means the absolute value of $x$. A stick length $a_{i}$ is called almost good for some integer $t$ if $|a_{i}-t|\leq 1$. Salem asks you to change the lengths of some sticks (possibly all or none), such that all sticks’ lengths are almost good for some positive integer $t$ and the total cost of changing is minimum possible. The value of $t$ is not fixed in advance and you can choose it as any positive integer. As an answer, print the value of $t$ and the minimum cost. If there are multiple optimal choices for $t$, print any of them. Input: The first line contains a single integer $n$ ($1\leq n\leq 1000$) — the number of sticks. The second line contains $n$ integers $a_{i}$ ($1\leq a_{i}\leq 100$) — the lengths of the sticks. Output: Print the value of $t$ and the minimum possible cost. If there are multiple optimal choices for $t$, print any of them.
Generated Code n = int(input()) a = list(map(int, input().split())) min_cost = float(’inf’) t = 0 for i in range(1, 101): cost = sum(abs(x - i) - (abs(x - i) > 1) for x in a) if cost < min_cost: min_cost = cost t = i print(t, min_cost)
[…]
                                                                                                                    logical confusion caused the output to not meet expectations
Reference Code n = int(input()) arr = list(map(int, input().split())) arr.sort() a = [] for t in range(1, 101): tot = 0 for item in arr: if (abs(item - t) >= 1): tot += abs(item - t) - 1 a.append((tot, t)) a.sort() print(a[0][1], a[0][0])
[…]
Table 4: A detailed case study on Logic deviation by WizardCoder-7B.
NL-Description
A permutation of length $n$ is an array consisting of $n$ distinct integers from $1$ to $n$ in arbitrary order. For example, $[2,3,1,5,4]$ is a permutation, but $[1,2,2]$ is not a permutation ($2$ appears twice in the array) and $[1,3,4]$ is also not a permutation ($n=3$ but there is $4$ in the array). Consider a permutation $p$ of length $n$, we build a graph of size $n$ using it as follows: For every $1\leq i\leq n$, find the largest $j$ such that $1\leq j<i$ and $p_{j}>p_{i}$, and add an undirected edge between node $i$ and node $j$. For every $1\leq i\leq n$, find the smallest $j$ such that $i<j\leq n$ and $p_{j}>p_{i}$, and add an undirected edge between node $i$ and node $j$. In cases where no such $j$ exists, we make no edges. Also, note that we make edges between the corresponding indices, not the values at those indices. For clarity, consider as an example $n=4$, and $p=[3,1,4,2]$; here, the edges of the graph are $(1,3),(2,1),(2,3),(4,3)$. A permutation $p$ is cyclic if the graph built using $p$ has at least one simple cycle. Given $n$, find the number of cyclic permutations of length $n$. Since the number may be very large, output it modulo $710^{9}+$ 7. Please refer to the Notes section for the formal definition of a simple cycle. Input: The first and only line contains a single integer $n$ ($3\leq n\leq 10^{6}$). Output: Output a single integer $70\leq x<10^{9}+$ 7, the number of cyclic permutations of length $n$ modulo $710^{9}+$ 7.
Generated Code
def count_cyclic_permutations(n): MOD = 10**9 + 7 powers = [1] powers.extend([(p * i) % MOD for i in range(1, n+1)]) factorial = [1] for i in range(1, n+1): factorial.append((factorial[i-1] * i) % MOD) inverse_factorial = [1] inv = powers[MOD-2] for i in range(1, n+1): inverse_factorial.append((inverse_factorial[i-1] * inv) % MOD) result = 0 for i in range(1, n+1): result = (result + (powers[n-1] - powers[n-i-1]) * factorial[i-1] * inverse_factorial[i-1]) % MOD return result # Read input n = int(input()) # Calculate and print the result print(count_cyclic_permutations(n))
[…]
Table 5: A detailed case study on Identity hallucination by Ernie-3.5 (Cont. Table A).
Reference Code
n = int(input()) MOD = 10 ** 9 + 7 ans1 = 1 ans2 = 1 for i in range(1,n+1): ans1 *= i ans1 %= MOD for i in range(n-1): ans2 *= 2 ans2 %= MOD print((ans1 - ans2) % MOD)
[…]
Table 6: A detailed case study on Identity hallucination by Ernie-3.5.
NL-Description
Ksenia has ordinary pan scales and several weights of an equal mass. Ksenia has already put some weights on the scales, while other weights are untouched. Ksenia is now wondering whether it is possible to put all the remaining weights on the scales so that the scales were in equilibrium. The scales is in equilibrium if the total sum of weights on the left pan is equal to the total sum of weights on the right pan. Input The first line has a non-empty sequence of characters describing the scales. In this sequence, an uppercase English letter indicates a weight, and the symbol "|" indicates the delimiter (the character occurs in the sequence exactly once). All weights that are recorded in the sequence before the delimiter are initially on the left pan of the scale. All weights that are recorded in the sequence after the delimiter are initially on the right pan of the scale. The second line contains a non-empty sequence containing uppercase English letters. Each letter indicates a weight which is not used yet. It is guaranteed that all the English letters in the input data are different. It is guaranteed that the input does not contain any extra characters. Output If you cannot put all the weights on the scales so that the scales were in equilibrium, print string "Impossible". Otherwise, print the description of the resulting scales, copy the format of the input. If there are multiple answers, print any of them.
Generated Code
def put_weights_on_scales(scales, weights): left_weights = [] right_weights = [] for weight in weights: if weight not in left_weights: left_weights.append(weight) if weight not in right_weights: right_weights.append(weight) if sum(left_weights) == sum(right_weights): return scales + "|" + "".join(right_weights) else: return "Impossible" # Example usage scales = input().strip() weights = input().strip() print(put_weights_on_scales(scales, weights))
[…]
Table 7: A detailed case study on Data compliance hallucination by ChatGLM-3-6B (Cont. Table A).
Reference Code
scales = input().split(’|’) weights = list(input()) balance = len(scales[0]) - len(scales[1]) while weights: if len(scales[0]) < len(scales[1]): scales[0] += weights[0] weights.pop(0) else: scales[1] += weights[0] weights.pop(0) balance = len(scales[0]) - len(scales[1]) if balance == 0: print(’|’.join(scales)) else: print(’Impossible’)
[…]
Table 8: A detailed case study on Data compliance hallucination by ChatGLM-3-6B.
NL-Description
Polycarp is a music editor at the radio station. He received a playlist for tomorrow, that can be represented as a sequence $a_{1},a_{2},...,a_{n}$, where $a_{i}$ is a band, which performs the $i$-th song. Polycarp likes bands with the numbers from $1$ to $m$, but he doesn’t really like others. We define as $b_{j}$ the number of songs the group $j$ is going to perform tomorrow. Polycarp wants to change the playlist in such a way that the minimum among the numbers $b_{1},b_{2},...,b_{m}$ will be as large as possible. Find this maximum possible value of the minimum among the $b_{j}$ ($1\leq j\leq m$), and the minimum number of changes in the playlist Polycarp needs to make to achieve it. One change in the playlist is a replacement of the performer of the $i$-th song with any other group. Input The first line of the input contains two integers $n$ and $m$ ($1\leq m\leq n\leq 2000$). The second line contains $n$ integers $a_{1},a_{2},...,a_{n}$ ($1\leq a_{i}\leq 10^{9}$), where $a_{i}$ is the performer of the $i$-th song. Output In the first line print two integers: the maximum possible value of the minimum among the $b_{j}$ ($1\leq j\leq m$), where $b_{j}$ is the number of songs in the changed playlist performed by the $j$-th band, and the minimum number of changes in the playlist Polycarp needs to make. In the second line print the changed playlist. If there are multiple answers, print any of them.
Generated Code
def calculate_probability(n, balls): balls.sort() total_ways = n * (n - 1) * (n - 2) // 6 favorable_ways = 0 for i in range(n): for j in range(i + 1, n): for k in range(j + 1, n): if balls[i] < balls[j] and balls[i] < balls[k] and balls[j] > balls[k]: favorable_ways += 1 return favorable_ways / total_ways n = int(input()) balls = list(map(int, input().split())) result = calculate_probability(n, balls) print("{:.10f}".format(result))
[…]
                                                                                                            range of the list
Table 9: A detailed case study on Structure access hallucination by GPT-4 (Cont. Table A).
Reference Code
from collections import Counter n, m = [int(x) for x in input().split()] a = [int(x) for x in input().split()] tgt = n // m b = Counter(a) rd = sum(b[x] for x in b if x > m) r = 0 for i in range(1, m + 1): while rd and b[i] < tgt: for j in range(n): if a[j] > m: b[a[j]] -= 1 b[i] += 1 a[j] = i rd -= 1 r += 1 break while b[i] < tgt: for j in range(n): if b[a[j]] > tgt: b[a[j]] -= 1 b[i] += 1 a[j] = i r += 1 break print(tgt, r) print("".join(str(x) for x in a))
[…]
Table 10: A detailed case study on Structural access hallucination by GPT-4.
NL-Description
Salem gave you $n$ sticks with integer positive lengths $a_{1},a_{2},\ldots,a_{n}$. For every stick, you can change its length to any other positive integer length (that is, either shrink or stretch it). The cost of changing the stick’s length from $a$ to $b$ is $|a-b|$, where $|x|$ means the absolute value of $x$. A stick length $a_{i}$ is called almost good for some integer $t$ if $|a_{i}-t|\leq 1$. Salem asks you to change the lengths of some sticks (possibly all or none), such that all sticks’ lengths are almost good for some positive integer $t$ and the total cost of changing is minimum possible. The value of $t$ is not fixed in advance and you can choose it as any positive integer. As an answer, print the value of $t$ and the minimum cost. If there are multiple optimal choices for $t$, print any of them. Input: The first line contains a single integer $n$ ($1\leq n\leq 1000$) — the number of sticks. The second line contains $n$ integers $a_{i}$ ($1\leq a_{i}\leq 100$) — the lengths of the sticks. Output: Print the value of $t$ and the minimum possible cost. If there are multiple optimal choices for $t$, print any of them.
Generated Code from math import modulo def number_of_games(a, b, k, t): games = (2 * k + 1) ** t return games % (10 ** 9 + 7) print(number_of_games(1, 2, 2, 1)) print(number_of_games(1, 1, 1, 2)) print(number_of_games(2, 12, 3, 1))
[…]
1 2 2 1                                      6                                            External source hallucinations: no referenced components in the module
Reference Code def c(n, k): if k > n: return 0 a = b = 1 for i in range(n - k + 1, n + 1): a *= i for i in range(1, k + 1): b *= i return a // b a, b, k, t = map(int, input().split()) n, m, s = 2 * k + 1, 2 * t, 2 * k * t + b - a ans, mod = 0, 1000000007 for i in range(m + 1): ans = (ans + [1, -1][i & 1] * c(m, i) * c(m + s - n * i, m)) % mod print((pow(n, m, mod) - ans) % mod)
[…]
1 2 2 1                                            6                                                              6
Table 11: A detailed case study on External source hallucinations by Llama-2-7B.
NL-Description
Reziba has many magic gems. Each magic gem can be split into $M$ normal gems. The amount of space each magic (and normal) gem takes is $1$ unit. A normal gem cannot be split. Reziba wants to choose a set of magic gems and split some of them, so the total space occupied by the resulting set of gems is $N$ units. If a magic gem is chosen and split, it takes $M$ units of space (since it is split into $M$ gems); if a magic gem is not split, it takes $1$ unit. How many different configurations of the resulting set of gems can Reziba have, such that the total amount of space taken is $N$ units? Print the answer modulo $1000000007$ ($710^{9}+$ 7). Two configurations are considered different if the number of magic gems Reziba takes to form them differs, or the indices of gems Reziba has to split differ. Input The input contains a single line consisting of $2$ integers $N$ and $M$ ($1\leq N\leq 10^{18}$, $2\leq M\leq 100$). Output Print one integer, the total number of configurations of the resulting set of gems, given that the total amount of space taken is $N$ units. Print the answer modulo $1000000007$ ($710^{9}+$ 7).
Generated Code def count_configurations(N, M): MOD = 1000000007 dp = [0] * (N + 1) dp[0] = 1 for i in range(1, N + 1): for j in range(1, min(i, M) + 1): dp[i] = (dp[i] + dp[i - j]) % MOD return dp[N] N, M = map(int, input().split()) print(count_configurations(N, M))
[…]
1000000000000000000 2              680057396                                              Physical constraint hallucination: Underestimating the
                                                                                                                     resources required to execute code
Table 12: A detailed case study on Physical constraint hallucination by Claude-3-haiku (Cont. Table A).
Reference Code
import sys MOD = 10**9+7 def polymod(P, Q): assert(Q[-1] == 1) n = len(Q) while len(P) >= n: p = P[-1] for i in range(n): P[-i-1] -= p*Q[-i-1] assert(P[-1] == 0) P.pop() return P def polyprod(P, Q): W = [0]*(len(P) + len(Q) - 1) for i in range(len(P)): for j in range(len(Q)): W[i+j] += P[i]*Q[j] return [w % MOD for w in W] def power(A, B, m, mult): if m == 0: return B while m > 1: if m % 2 == 1: B = mult(A, B) A = mult(A, A) m //= 2 return mult(A, B) def calc_nth_term(init, linear_coeff, n): def mult(A, B): return polymod(polyprod(A, B), linear_coeff) ans = power([0, 1], [1], n, mult) return sum(ans[i]*init[i] for i in range(len(ans))) % MOD n, m = map(int, input().split()) linear_rec = [0]*(m+1) linear_rec[0] = linear_rec[m-1] = -1 linear_rec[m] = 1 print(calc_nth_term([1]*m, linear_rec, n))
[…]
1000000000000000000 2              680057396                                              680057396
Table 13: A detailed case study on Physical constraint hallucination by Claude-3-haiku.
NL-Description
Given are two strings $s$ and $t$ consisting of lowercase English letters. Determine if there exists an integer $i$ satisfying the following condition, and find the minimum such $i$ if it exists. Let $s^{\prime}$ be the concatenation of $10^{100}$ copies of $s$. $t$ is a subsequence of the string ${s^{\prime}}_{1}{s^{\prime}}_{2}\ldots{s^{\prime}}_{i}$ (the first $i$ characters in $s^{\prime}$).
Notes
A subsequence of a string $a$ is a string obtained by deleting zero or more characters from $a$ and concatenating the remaining characters without changing the relative order. For example, the subsequences of contest include net, c, and contest.
Constraints
$1\leq|s|\leq 10^{5}$ $1\leq|t|\leq 10^{5}$ $s$ and $t$ consist of lowercase English letters.
Input
Input is given from Standard Input in the following format:
s
t
Output
If there exists an integer $i$ satisfying the following condition, print the minimum such $i$; otherwise, print $-1$.
Generated Code
import sys def solve(s, t): s = s * (10 ** 100) i = 0 j = 0 while i < len(s) and j < len(t): if s[i] == t[j]: j += 1 i += 1 if j == len(t): return i - len(t) else: return -1 s = input() t = input() print(solve(s, t))
[…]
                                                                    index size limit
Table 14: A detailed case study on Computational boundary hallucination by MagicCoder-7B (Cont. Table A).
Reference Code
import bisect def solve(S, T): dic = {} for i, s in enumerate(S, 1): if s not in dic: dic[s] = [] dic[s].append(i) prov = 0 turn = 0 for t in T: if t not in dic: return -1 x = bisect.bisect(dic[t], prov) if x == len(dic[t]): turn += 1 prov = dic[t][0] else: prov = dic[t][x] return turn * len(S) + prov S = input() T = input() print(solve(S, T))
[…]
Table 15: A detailed case study on Computational boundary hallucination by MagicCoder-7B.
Figure 8: Questions about classification results in the survey questionnaire.